BenMabe:

A young sumo wrestler decided to go on a special high-protein diet to gain weight rapidly. After 1 months, he weighed 140 kilograms. He gained weight at a rate of 5.5 kilograms per month. Let y represent the sumo wrestler's weight (in kilograms) after x months.

TheSmartOne:

So how would you go about this equation? Hint: Current weight = (rate at which he's increasing)*(some variable for months) + (what he originally weighs)

TheSmartOne:

it'll be in a similar y = mx + b format

BenMabe:

ok that seems easy

TheSmartOne:

It is You just have to calculate his original weight and everything else is pretty much given Let me know what you get!

BenMabe:

im either getting 145.5 or 140.5

BenMabe:

and im also simple minded so this seems easy but for me when i think about it its hard

TheSmartOne:

In one month he gained 5.5 kg So he now weighs 140 kg How much did he weigh before he gained 5.5 kgs?

BenMabe:

135 kg right

BenMabe:

134.5 kg

TheSmartOne:

Becareful it's 5.5 not 5

TheSmartOne:

Spot on

BenMabe:

ok so before he did this diet he weighted 134.5 and then 1 month later he weights 140

BenMabe:

so what then

BenMabe:

the 1 month is meant to be 11 my bad heh

TheSmartOne:

Yes So we have so the information we need Look at that original hint I gave you y = mx + b y and x are our variables (they're just random letters where we can plug stuff into one of them to get a number for the other letter) x is for months y is for his final weight after x months m is our slope or you can call it the rate of change And the question gives us this. He's gaining 5.5 kg every month so our m is 5.5 B is the y-intercept. This is the weight he started off with and we just calculated it. So can you put it so together now?

BenMabe:

140=5.5*11+134.5

TheSmartOne:

Hmmm we want an equation so that means we need to keep the x and y as is So that way we can plug in a number for the month and find out how much he weighs now (Also I think you meant you 1 and not 11 because otherwise that math wouldn't work out haha)

BenMabe:

well i made a mistake and i didnt know it but when i looked back a my question it said 11 so idk sry

BenMabe:

Y=5.5x+134.5 is that how u would put it

TheSmartOne:

If you put x and y back in to it, it'll all be good

TheSmartOne:

yay! good job!! You're getting the hang of it

BenMabe:

ok so let get this right the right answer is going to be Y=5.5x+134.5 correct

TheSmartOne:

Yup!

BenMabe:

ok thank you

TheSmartOne:

No problem!